OpenStudy (anonymous):

find the value of x that makes the triangles similar:

Directrix (directrix):

Lengths of correponding sides of similar triangles have to be in proportion. Side with lengths 28 and 42 are corresponding which makes the scale factor 28/42 or 2/3. 2/3 = 22/(5x + 3) Cross multiply and solve for x.

Directrix (directrix):

@angelaaaxo

OpenStudy (anonymous):

oh so you have to make sure the triangles are both the right way like those are and once they are you just take one side and do the copy over orginial?

OpenStudy (anonymous):

idk im so confused :S

OpenStudy (anonymous):

i did 28/42 = 2/3 2/3 = 22/5x+3 like you said so would it be 10x+6 ?lol

Directrix (directrix):

Solve for x 28/42 = 22/(5x + 3) That forces 22 and (5x + 3) to have the same ratio as 28 over 42.

OpenStudy (anonymous):

thank you so much :D

OpenStudy (anonymous):

K

OpenStudy (anonymous):

@Luis_Rivera can u help me on a problem??